OpenStudy (anonymous):

HELP WITH ANSWER PLEASE , ILL MEDAL Doris has a savings account at her local bank. She started with $500 and deposits $150 every month. The equation below shows her balance, b, after m months. What is the balance in Doris' account after 4 months? b = 150m + 500 A. $600 B. $650 C. $800 D. $1100

OpenStudy (amistre64):

let m=4, what is b?

OpenStudy (anonymous):

Um , is it b=600 ? or 650 ? @amistre64

OpenStudy (amistre64):

150(4) = 150+150+150+150 = 400+200, then we add in 500 so im thinking its gonna be more that 650